1.Put the equation 3x - 2y = 16 in slope-intercept form.
A.y = -3x + 16
B.-y = 3x - 8
C.y = -5x + 16
D.y = (3/2)x - 8


2.If y varies inversely as x, and y = 8 when x = 4, find x when y = 2
A.1
B.16
C.4
D.8




Answer by SandySharma(35)

1.Put the equation 3x - 2y = 16 in slope-intercept form.
3x - 2y = 16
Subtract 3x from both sides, you will get
-2y = -3x + 16
Now, divide both sides by -2
y = -3x/-2 + 16/-2
y = 3x/2 - 8
So, answer is D.



2.If y varies inversely as x, and y = 8 when x = 4, find x when y = 2
A.1
B.16
C.4
D.8
Given that, y varies inversely as x
that is y = k/x
y = 8 when x = 4
substitute in the equation
8 = k/4
we will get, k = 32
Now find x when y = 2 & k= 32,
2 = 32/x
x = 16
So, answer is B.